40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jul 8 15:31:48 2013, Jenne, Update, ASC, POP QPD calibration prep 
    Reply  Tue Jul 9 11:41:22 2013, Jenne, Update, ASC, POP QPD calibration attempt 
       Reply  Tue Jul 9 16:08:32 2013, Jenne, Update, ASC, POP QPD calibration attempt 
          Reply  Fri Jul 12 21:23:42 2013, Jenne, Update, ASC, POP QPD calibration attempt ChangeVertMicrometer_July2013.pngChangeSideMicrometer_July2013.png
             Reply  Mon Jul 15 17:16:59 2013, Jenne, Update, ASC, POP QPD calibration attempt ChangeSideMicrometer_July2013_calib2.pngChangeSideMicrometer_July2013_calib1.pngChangeVertMicrometer_July2013_calib1.png
                Reply  Tue Jul 16 01:17:21 2013, Jenne, Update, ASC, POP QPD calibration attempt 
                   Reply  Tue Jul 23 01:30:27 2013, Jenne, Update, ASC, POP QPD analysis 9x
                      Reply  Fri Jul 26 13:39:30 2013, Koji, Update, ASC, POP QPD analysis 
Message ID: 8812     Entry time: Tue Jul 9 16:08:32 2013     In reply to: 8810     Reply to this: 8840
Author: Jenne 
Type: Update 
Category: ASC 
Subject: POP QPD calibration attempt 

[Jenne, Alex] 

Calibration data for the POP QPD has been taken, with the PRMI locked on sideband (with AS55Q and REFL33I, since it stayed locked longer with those 2).  ASC was on, AC coupled. 

We didn't get too far on either side of center of the QPD, since the ASC servo would go unstable, so we only explored the roughly linear region.  Data / plots / analysis to follow.

ELOG V3.1.3-